OpenStudy (anonymous):

Sketch the graph of f and use it to find the domain, range, and to determine whether the function is one-to-one. f(x)=root16-x^2

OpenStudy (amistre64):

\[\sqrt{16-x^2}\] is this the equation?

OpenStudy (anonymous):

Yes.

OpenStudy (amistre64):

the graph looks like angel wings.... going off in both directions

OpenStudy (amistre64):

its not a one to one function

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=sqrt%2816-x^2%29 lol... ok, not angel wings, but still not one to one

OpenStudy (anonymous):

it is the upper half of a circle with radius 4

OpenStudy (anonymous):

the equation for a circle centered at the origin with radius 4 is \[x^2+y^2=4^2\]

OpenStudy (anonymous):

Yeah, that's what I have on my calculator. I don't remember how to do the rest though....

OpenStudy (anonymous):

if you solve this equation for y you get \[y=\pm\sqrt{1-x^2}\]

OpenStudy (anonymous):

excuse me i mean \[y=\pm\sqrt{16-x^2}\]

OpenStudy (anonymous):

if you only take the positive part you are getting just the stuff above the x axis.

OpenStudy (anonymous):

it is in fact a one to one funciton

OpenStudy (anonymous):

Okay, that makes sense why it only shows up half on my graph.

OpenStudy (anonymous):

oops no it isn't sorry

OpenStudy (anonymous):

doesn't pass horizontal line test.

OpenStudy (amistre64):

when x = -4 OR +4 we get the same y

OpenStudy (anonymous):

domain is [-4,4] range is [0,4]

OpenStudy (anonymous):

of course i was being stupid

myininaya (myininaya):

satellite!

OpenStudy (anonymous):

55 to go i am in awe. but might have to wait until tomorrow yes?

OpenStudy (anonymous):

hello myininaya how goes it?

OpenStudy (amistre64):

maybe, with the holidays its hard to get good clickers lol

OpenStudy (anonymous):

This is true, lol. Thanks guys<333

myininaya (myininaya):

i need help this guy is totally confusing me about stuff i dont what the frick hes talking about

OpenStudy (anonymous):

yeah really. who does math on memorial day?

OpenStudy (anonymous):

what question?

OpenStudy (anonymous):

I do, apparently :P

myininaya (myininaya):

follow me